Project

General

Profile

Actions

1. Commissioning and Ops Scheduling meeting : 17/11/2015

To be held in CR10 in the ATLAS Building @ 15:30 pm

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Agenda:

  • Review of tasks and activities identified at last meeting
    • Focus coil flip mode commissioning
    • SS Power supply and vacuum work
    • Absorber leak testing and cooldown plans
    • Diffuser leak fix
    • Decay Solenoid PSU Repair
    • Result of discussion on compressed air transducer installation
  • Upcoming activities : chaired by Victoria
    • Work plan
    • Control room systems test (Wed, 18th)
    • Absorber and SSD leak testing (Wed, 18th)
    • Resurvey of TOF1 (23rd)
    • Expert-led run (24th)
    • Hydrogen/Vacuum systems installation and commissioning (Phi Warburton, Tue-Wed, 24th--25th)
    • Conventional magnet cooling system flush (25th)
    • Focus coil flip mode commissioning (26th)
    • Others
  • AOB

Minutes

  • Review
    • Focus coil flip mode commissioning : happened on Thursday. Everything looked OK up to 214 A at which point FC quenched. Another soak test scheduled for Nov 26th. May need to make the ramping more stable by doing is more slowly. Next soak test might involve a longer ramp with rests at various currents. John Cobb suggested that we should plan for at least 2 more tests since getting to 225 A might be tricky.
    • SS power supply and vacuum : more work to do extending into next week.
    • Absorber leak testing and cool down : last Thursday absorber started cooling down and then bottomed off. Now warming and Steve wants to do some leak testing this week.
    • Diffuser leak fix : Under way now. Diffuser removed easily and blanking plate installed. Diffuser in Oxford and will be fitted into new
      top hat. Back on Thursday for re-installation and testing.
    • Decay Solenoid PSU repair : PSU will be removed on Wednesday 18th. Need a decision on whether the PSU can be fixed on-site or sent back to FUG in Bavaria. FUG have been consulted on the PS. Henry suggests taking it out and putting in R9. Mike says that it can be fixed in R9 but not tested.
      Mike wants the right people to be involved in the discussion : Mark Tucker, Josef, himself, the Daresbury team co-ordinated by Colin. The movement and repair are decoupled so we need to decide what to do with it now. Mike will try to resolve this issue.
    • Transducer issue : discussion happened and a transducer is not needed.
  • Plan
    • Work plan to Friday online
    • Search the Hall on Monday evening for expert run on Tuesday. This clashes with Phil Warburton's work on Tuesday/Wednesday. Need
      to understand if Phil only needs Wednesday? Test run still nominally 24th but need to talk to Phil.
    • Conventional magnet Q4 cooling system flush will take place on Wednesday as there is access to the DSA. Schedule Wednesday afternoon to test conventional magnets after work.
    • Repeat focus coil commissioning on Thursday, possibly for longer to allow stabilisation of the FC ramp.
    • Can't currently leave the FC powered overnight with no monitoring. We need to discuss how we allow this, ask ISIS to do the monitoring or some other procedure before we leave the FC overnight.
  • AOB
    • None.

Updated by Boyd, Steven over 7 years ago ยท 9 revisions